# Meta-automatic Sequences

### v1

@@ 5.2 Meta(1) and Meta(2)
The incidence matrices of the morphisms in Corollaries 25 and 32 are
$M_{\mathcal{M}_{1}}=\begin{pmatrix}1&2&1&0\\ 1&1&1&1\\ 1&1&1&1\\ 1&0&1&2\end{pmatrix},\qquad M_{\mathcal{M}_{2}}=\begin{pmatrix}1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\end{pmatrix}$.
One checks $M_{\mathcal{M}_{1}}^{2}>0$ and $M_{\mathcal{M}_{2}}>0$ directly, so both morphisms are primitive. From the work of Pansiot [24], the fixed point of a primitive $k$-uniform morphism over an alphabet $\mathcal{A}$ has complexity $p(n)=\Theta(n)$. Each of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is a coding $\tau(\sigma^{\omega}(a_{0}))$, where $\tau\colon\mathcal{A}\to\{0,1\}$ is non-eventually-constant. Since every letter appears with positive frequency and $\tau$ takes both values, distinct factors of $\sigma^{\omega}(a_{0})$ not identified by $\tau$ remain distinct in $\tau(\sigma^{\omega}(a_{0}))$, giving $p_{\tau(\sigma^{\omega})}(n)\geq p_{\sigma^{\omega}}(n)/|\mathcal{A}|^{2}=\Theta(n)$ [3, Proposition 10.9.2]. ∎
The exact complexity at dyadic lengths, and the complete piecewise formula, are given in the theorems below.
@@
In particular, the dyadic slope alternates between $\frac{5}{2}$ and $3$ according to $k\bmod 3$. The three cases modulo $3$ are tied together by the recurrence
$p_{\mathcal{M}_{2}}(2^{k+3})=8\,p_{\mathcal{M}_{2}}(2^{k})+14\qquad(k\geq 3).$
@@
The formula is certifiable by the Walnut theorem prover, leaving the details to the reader. ∎
Since $\mathrm{RS}_{\mathcal{M}_{2}}(n)\in\{2,4\}$, the values of $p_{\mathcal{M}_{2}}$ between dyadic points are determined by the support of $\mathrm{RS}_{\mathcal{M}_{2}}=4$. This support has a precise structure that extends Theorem 45 to a complete evaluation.
@@
For $n\geq 1$, set $k=\lfloor\log_{2}n\rfloor$ and $r=k\bmod 3$.
Case $r=2$: On the entire block $n\in[2^{k},2^{k+1}]$,
@@
Case $r=0$: The block $n\in[2^{k},2^{k+1}]$ splits into three linear pieces,
@@
Case $r=1$: The block $n\in[2^{k},2^{k+1}]$ splits into three linear pieces,
@@
Again, a proof certificate can be obtained via Walnut, leaving the details to the reader. ∎
6 Concluding remarks
A parallel direction replaces the dyadic balancedness given by $a(2n)+a(2n+1)=1$ by the ternary constraint $a(3n)+a(3n+1)+a(3n+2)=2$, while still restricting to binary values in $\{0,1\}$. This forces exactly two $1$s and one 0 in each base-$3$ block, giving density $2/3$ on prefixes of length $3^{m}$. Meta-automatic selectors of the form a(3n+ra(n))a(3n+r-a(n)) arise naturally, and the $\mathbb{F}_{2}$-linearization mechanism of Theorem 23 extends to this setting: while such selectors are not denestable, the ternary balance constraint may again convert them into fixed-argument $\mathbb{F}_{2}$-affine maps, yielding $3$-automatic sequences. A small number of initial configurations appear to produce binary $3$-automatic sequences with small DFAOs; exploring this family and its complexity functions is left for a separate treatment.
@@ 7 Appendix
Table 1: Minimal DFAO size for each specialization. “Sel.” counts the number of meta-automatic selectors among $F$ and $G$. Boldface marks the three sequences studied in this paper.
$F(n)$ $G(n)$ DFAO Sel. Notes
$a(n)$ $a(n)$ 22 0 Thue–Morse
$a(n)$ $a(2n)$ $\mathbf{3}$ $\mathbf{0}$ Quarto $\mathcal{Q}$
$a(n)$ $a(2n+1)$ 22 0 Thue–Morse
$a(n)$ $a(2n+a(n))$ 44 11
$a(n)$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{1}$ Meta(1) $\mathcal{M}_{1}$
$a(2n)$ $a(n)$ 44 0
@@
$a(2n)$ $a(2n+1)$ 22 0 Thue–Morse
$a(2n)$ $a(2n+a(n))$ 33 11
$a(2n)$ a(2n+1a(n))a(2n+1-a(n)) 33 11
$a(2n+1)$ $a(n)$ 55 0
$a(2n+1)$ $a(2n)$ 33 0 complement of $\mathcal{Q}$
$a(2n+1)$ $a(2n+1)$ 55 0
$a(2n+1)$ $a(2n+a(n))$ 55 11
$a(2n+1)$ a(2n+1a(n))a(2n+1-a(n)) 55 11
$a(2n+a(n))$ $a(n)$ 44 11
$a(2n+a(n))$ $a(2n)$ 44 11
$a(2n+a(n))$ $a(2n+1)$ 44 11
$a(2n+a(n))$ $a(2n+a(n))$ 44 22
$a(2n+a(n))$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{2}$ Meta(2) $\mathcal{M}_{2}$
a(2n+1a(n))a(2n+1-a(n)) $a(n)$ 44 11
a(2n+1a(n))a(2n+1-a(n)) $a(2n)$ 44 11
a(2n+1a(n))a(2n+1-a(n)) $a(2n+1)$ 44 11
a(2n+1a(n))a(2n+1-a(n)) $a(2n+a(n))$ 44 22 complement of $\mathcal{M}_{2}$
a(2n+1a(n))a(2n+1-a(n)) a(2n+1a(n))a(2n+1-a(n)) 44 22
@@ 8 Acknowledgments
Claude Opus 4.6 (Anthropic) was used as a computational assistant, in particular for exploring the factor-complexity functions of $\mathcal{Q}$, $\mathcal{M}_{2}$, and $\mathcal{M}_{1}$ of Section 5 and guessing the formulas, and systematically investigating variants of the recursions as in Section 7, which allowed us to select $\mathcal{Q}$, $\mathcal{M}_{1}$, and $\mathcal{M}_{2}$ as three sequences of particular interest for our purposes. The first author thanks Jean-Paul Allouche, Karl Dilcher, and Jeffrey Shallit for useful feedback related to this paper.

### v2

@@ 5.2 Factor complexity of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$
The incidence matrices of the morphisms in Corollaries 16 and 23 are
$M_{\mathcal{M}_{1}}=\begin{pmatrix}1&2&1&0\\ 1&1&1&1\\ 1&1&1&1\\ 1&0&1&2\end{pmatrix},\qquad M_{\mathcal{M}_{2}}=\begin{pmatrix}1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\\ 1&1&1&1\end{pmatrix}$. (19)
The incidence matrices in (19) give us that both of the associated morphisms are primitive. A result of Pansiot [26] gives us that the fixed point of a primitive $k$-uniform morphism has factor complexity $O(n)$, and the same upper bound holds after applying a coding. Neither $\mathcal{M}_{1}$ nor $\mathcal{M}_{2}$ is ultimately periodic, as can be checked from the DFAOs of Theorems 13 and 22, or certified by Walnut. By the Morse–Hedlund theorem [3, Theorem 10.2.6], the factor complexity of a non-ultimately-periodic word is at least $n+1$ for every $n\geq 0$. Hence both complexities are $\Theta(n)$. ∎
An explicit evaluation of the factor complexity of $\mathcal{M}_{2}$ is provided below.
@@
Moreover, the recurrence
$p_{\mathcal{M}_{2}}(2^{k+3})=8\,p_{\mathcal{M}_{2}}(2^{k})+14\qquad(k\geq 3)$
holds true.
@@
This is certified by Walnut. ∎
Since $\mathrm{RS}_{\mathcal{M}_{2}}(n)\in\{2,4\}$, the values of $p_{\mathcal{M}_{2}}$ are determined by the positions at which $\mathrm{RS}_{\mathcal{M}_{2}}(n)=4$. The following proposition extends Proposition 36 to a complete evaluation.
@@
We have $p_{\mathcal{M}_{2}}(1)=2$. For $n\geq 2$, set $k=\lfloor\log_{2}n\rfloor$ and $r=k\bmod 3$.
Case $r=2$: On the interval $2^{k},2^{k+1}]$, we have that
@@
Case $r=0$: On the interval $2^{k},2^{k+1}]$, we have that
@@
Case $r=1$: On the interval $2^{k},2^{k+1}]$, we have that
@@
This is certified by Walnut. ∎
6 Conclusion
Consider replacing $a(2n)+a(2n+1)=1$ by $a(3n)+a(3n+1)+a(3n+2)=2$, with $a$ taking values in $\{0,1\}$. This gives rise to expressions of the form a(3n+ra(n))a(3n+r-a(n)), and the analogue of Lemma 5 converts each such expression into a fixed-argument expression. A small number of initial configurations produce $3$-automatic binary sequences with small DFAOs; we plan to treat this family separately.
@@ 7 Appendix
Table 1: Minimal DFAO size for each specialization. The column “Nested” counts the number of branch rules ($F$ and $G$) whose right-hand side contains an expression of the form $a(\,\cdot\pm a(n))$. Boldface marks the three sequences studied in this paper.
$F(n)$ $G(n)$ DFAO Nested Notes
$a(n)$ $a(n)$ 2 0 Thue–Morse
$a(n)$ $a(2n)$ $\mathbf{3}$ $\mathbf{0}$ $\mathcal{Q}$ (Thue–Morse Quarto)
$a(n)$ $a(2n+1)$ 2 0 Thue–Morse
$a(n)$ $a(2n+a(n))$ 4 1
$a(n)$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{1}$ $\mathcal{M}_{1}$
$a(2n)$ $a(n)$ 4 0
@@
$a(2n)$ $a(2n+1)$ 2 0 Thue–Morse
$a(2n)$ $a(2n+a(n))$ 3 1
$a(2n)$ a(2n+1a(n))a(2n+1-a(n)) 3 1
$a(2n+1)$ $a(n)$ 5 0
$a(2n+1)$ $a(2n)$ 3 0 complement of $\mathcal{Q}$
$a(2n+1)$ $a(2n+1)$ 5 0
$a(2n+1)$ $a(2n+a(n))$ 5 1
$a(2n+1)$ a(2n+1a(n))a(2n+1-a(n)) 5 1
$a(2n+a(n))$ $a(n)$ 4 1
$a(2n+a(n))$ $a(2n)$ 4 1
$a(2n+a(n))$ $a(2n+1)$ 4 1
$a(2n+a(n))$ $a(2n+a(n))$ 4 2
$a(2n+a(n))$ a(2n+1a(n))a(2n+1-a(n)) $\mathbf{4}$ $\mathbf{2}$ $\mathcal{M}_{2}$
a(2n+1a(n))a(2n+1-a(n)) $a(n)$ 5 1
a(2n+1a(n))a(2n+1-a(n)) $a(2n)$ 4 1
a(2n+1a(n))a(2n+1-a(n)) $a(2n+1)$ 5 1
a(2n+1a(n))a(2n+1-a(n)) $a(2n+a(n))$ 5 2
a(2n+1a(n))a(2n+1-a(n)) a(2n+1a(n))a(2n+1-a(n)) 5 2
@@ 8 Acknowledgments
The authors thank the anonymous referee for a careful reading of the manuscript and for helpful suggestions. The authors also thank Gandhar Joshi for independently testing the Walnut files. Claude (Anthropic) was used as a computational assistant, in particular for exploratory computations, for suggesting candidate formulas later checked by independent scripts, and for checking consistency between the manuscript, the Python script, and the Walnut files. The mathematical content, the proofs, and the final wording are the authors’ own. The first author thanks Jean-Paul Allouche, Karl Dilcher, and Jeffrey Shallit for useful feedback related to this paper.